# Empirical distribution of ancestral lineages in populations with density-dependent interactions

### v2

@@ 1 Introduction
In order to address this limitation, there have been developments towards capturing the ancestral lineage of a sampled individual, as well as the whole genealogical tree, in populations with interactions. Recently, a spinal construction has been developed for this setting, focusing on multi-type processes with discrete type space [4]. The general idea consists in biasing the reproduction rates of the process, both along the spine and outside of it, according to a positive function $\psi$ of the reproducing particle’s trait $x$ and the population’s type composition $\mathbf{z}$. Intuitively, $\psi(x,\mathbf{z})$ can be regarded as the individual’s reproductive value or long-term fertility. Hence, when the spine reproduces, descendances with higher values of $\psi$ given the population state are favored, while the descendances of individuals outside of the spine are biased towards rendering the population more favorable for the spine. This spinal construction has since been extended to include more general type spaces [23]. It further has served to study the convergence of genealogies of density-dependent branching processes to the Kingman coalescent, establishing a connection with backward-in-time approaches [1].
@@
In this paper, we focus on the empirical type distribution of ancestral lineages ($\psi=1$), which arguably corresponds to the most natural and naive sampling strategy. In particular, we aim to capture the underlying survivorship bias. Notably, in the many-to-one formula associated to the aforementioned $\psi$-spine, spinal trajectories are penalized by an exponential weight which is generally stochastic. Thus the interpretation of the survivorship bias embodied by this spinal process is not straight-forward, as this penalization needs to be taken into account. In addition, Monte-Carlo estimations of the many-to-one formula become delicate as rare trajectories may have a tremendous impact. As a consequence, we aim at proposing an alternative spinal construction, whose associated many-to-one formula does not require exponential weighting of trajectories. This is achieved by a time-inhomogeneous spinal process, inspired from a similar auxiliary process which captures the lineage of a uniformly sampled individual in a branching process with large initial population [22].
@@
Further, a natural regime to consider is the large population limit $K\to+\infty$. Under classical assumptions, several large population approximations can frequently be established. Indeed, these approximations may be more prone to mathematical analysis than the original population process. From a numerical point of view, simulating large population approximations is often less expensive than individual-based models for large population sizes, both in terms of computation time and memory. For instance, numerically evaluation of the $\psi$-spine many-to-one formula requires simulations of the spinal population process, which is computationally expensive for large population sizes. As a consequence, it is natural to develop approximate sampling strategies, which allow to sample directly in those large population approximations. We establish a many-to-one formula for sampling in the deterministic large population limit with our time-inhomogeneous spinal process. Finally, we quantify the approximation error which is committed by sampling in the large population limit instead of the original population process.
@@ 3 Survivorship bias and the empirical distribution of ancestral lineages
In this section, our aim is to propose a many-to-one formula without stochastic penalization of spinal trajectories, in an effort to gain insight on the survivorship bias. In order to do so, we will introduce a time inhomogeneous spinal construction. This approach is inspired by [22]. While we are particularly interested in the empirical distribution of ancestral lineages, the results obtained in this Section hold for a large class of sampling strategies, and are thus stated in a general setting.
@@ 3.1 A many-to-one formula
Let us now introduce the time-inhomogeneous spinal process, which allows to capture the behavior of the ancestral lineage of a $\psi$-weighted sample of the population process. For $t\geq 0$ fixed, we will consider the time-inhomogeneous Markov process $(Y^{(t)}(s),\zeta^{(t)}(s))_{s\leq t}$ defined as follows. The main idea is to follow the type $Y^{(t)}$ of a distinguished individual, which will be referred to as spine in analogy to classical spinal constructions. At time $s\leq t$, when of type $x$ in a population of state $\mathbf{z}$, the spine divides to leave descendance $\mathbf{k}$ and switch to type $y$ with rate
@@
In other words, compared to the original process, at any time $s\leq t$, transitions along the distinguished lineage are biased in favor of those which lead to a larger $\psi$-average descendance at the final time $t$. However, due to the density-dependence of division rates, it is necessary to keep track of the population state $\zeta^{(t)}$. Again, transitions need to be biased, in order to account for the modified behavior of the distinguished individual when compared to the original process. As a consequence, when the population is in state $\mathbf{z}$ and the spine of type $x$ at time $s\leq t$, individuals of type $y$ other than the spine reproduces to leave descendance $\mathbf{k}$ at rate
@@
Here, the bias favors those transitions which lead to a more favorable environment for the spine, i.e. a population composition in which the $\psi$-average of the spine’s descendance is high.
@@
$Y^{(t)}$ $(s)=Y^{(t)}(0)+\int_{0}^{s}\int_{E}\mathbf{1}_{\left\{\theta\leq\rho^{(t)}_{y,\mathbf{k}}(r,Y^{(t)}(r-),\zeta^{(t)}(r-))\right\}}(\mathbf{e}(y)-Y^{(t)}(r-))Q(dr,d\theta,n(dy,d\mathbf{k})),$ (5)
$\zeta^{(t)}$ $(s)=\zeta^{(t)}(0)+\int_{0}^{s}\int_{E}\mathbf{1}_{\left\{\theta\leq\rho^{(t)}_{y,\mathbf{k}}(r,Y^{(t)}(r-),\zeta^{(t)}(r-))\right\}}(\mathbf{k}-Y^{(t)}(r-))Q(dr,d\theta,n(dy,d\mathbf{k}))$
$+\int_{0}^{s}\int_{E}\mathbf{1}_{\left\{\theta\leq(\zeta^{(t)}_{y}(r-)-\mathbf{1}_{\left\{Y^{(t)}(r-)=y\right\}})\widehat{\rho}^{(t)}_{\mathbf{k}}(r,y,Y^{(t)}(r-),\zeta^{(t)}(r-))\right\}}(\mathbf{k}-\mathbf{e}(y))\widehat{Q}(dr,d\theta,n(dy,d\mathbf{k})).$
@@
This result shows that the reproduction rates defined in Equations (3) and (4) yield the appropriate survivorship bias: reproduction events are favored if and only if they increase the $\psi$-average of the descendance at sampling time. In particular, the survivorship bias may thus be computed explicitly, giving access to a more precise interpretation. This will be illustrated in Section 3.3.
@@ 3.2.2 Proof of the many-to-one formula
Let us now turn to the inhomogeneous semi-group property. Let $r\leq\tau\leq t$, and consider $f:\mathcal{S}_{K}\to\mathbb{R}_{+}$ and $(x,\mathbf{z})\in\mathcal{S}_{K}$. Throughout the proof, we let $\mathfrak{X}_{0}=\mathfrak{X}(\mathbf{z})$. By definition of the semi-group,
@@
$\mathbb{E}[\sum_{\begin{subarray}{c}u\in\mathbb{G}(t)\\ u\succeq u_{x}(r)\end{subarray}}\psi(x_{u}(t),Z(t))m(x_{u}(\tau),Z(\tau),t-\tau)^{-1}G((x_{u}(s),Z(s))_{s\leq\tau}))|X(r)=\mathfrak{X}_{0}]$ (15)
$=\mathbb{E}[\!\!\sum_{\begin{subarray}{c}v\in\mathbb{G}(\tau)\\ v\succeq u_{x}(r)\end{subarray}}\!\!\mathbb{E}[\!\!\sum_{\begin{subarray}{c}u\in\mathbb{G}(t)\\ u\succeq v\end{subarray}}\!\!\psi(x_{u}(t),Z(t))|X(\tau)]m(x_{v}(\tau),Z(\tau),t-\tau)^{-1}G((x_{v}(s),Z(s))_{s\leq\tau}))|X(r)=\mathfrak{X}_{0}]$
$=\mathbb{E}[\sum_{\begin{subarray}{c}v\in\mathbb{G}(\tau)\\ v\succeq u_{x}(r)\end{subarray}}G((x_{v}(s),Z(s))_{s\leq\tau}))|X(r)=\mathfrak{X}_{0}].$
@@ 3.3 An application to the empirical distribution of ancestral lineages in a density-dependent population
The particles behave as follows. If the population corresponds to a single particle of type $A$, the latter may give birth to another particle of the same type at rate $b$. Whenever there are two particles of type $A$, each may die due to competition at rate $d$, or escape competition by switching to state $B$ at rate $c_{A}$. Finally, a particle of type $B$ can only switch back to state $A$, at rate $c_{B}$. Throughout the following, we assume that $b=d$ and $c_{A}=c_{B}$, for ease of computation. In other words, the state space of the population with a distinguished particle is described by the ordered set $\{(A,1,0),(A,2,0),(A,1,1),(B,1,1)\}$, and the dynamics are characterized by four reproduction rates :
@@
$m(t)=\frac{4}{5}\begin{pmatrix}2\\ 1\\ 1\\ 1\\ \end{pmatrix}-\frac{\lambda_{-}}{10\sqrt{\Delta}}\mathrm{e}^{-\frac{\lambda_{+}t}{2}}\begin{pmatrix}-\frac{-3b+3c-\sqrt{\Delta}}{2c}\\ -\frac{-3b+c+\sqrt{\Delta}}{2m}\\ 1\\ 1\\ \end{pmatrix}+\frac{\lambda_{+}}{10\sqrt{\Delta}}\mathrm{e}^{-\frac{\lambda_{-}t}{2}}\begin{pmatrix}-\frac{-3b+3c+\sqrt{\Delta}}{2m}\\ -\frac{-3b+c-\sqrt{\Delta}}{2c}\\ 1\\ 1\\ \end{pmatrix}$,
@@ 4 Empirical distribution of ancestral lineages in large populations
In this section, we are interested in empirical distribution of ancestral lineages in large populations, corresponding to the case $\psi=1$. Indeed, when the population size $K$ grows to infinity, the population trait distribution converges under classical assumptions to a deterministic limit $z$, characterized as the unique solution of a dynamical system. Here, we show that it is possible to sample in the limit $z$ with a time-inhomogeneous spine construction, analogously to Theorem 3.3. In addition, we quantify the approximation error committed by sampling in the large population limit instead of the original finite-size population process.
@@ 4.1 Many-to-one formula and speed of convergence
For $K\geq 1$, consider the population process $X^{K}$ where an individual of type $x$ in a population of composition $\mathbf{z}\in\mathcal{Z}_{K}$ is replaced by descendance $\mathbf{k}$ at rate $\tau_{\mathbf{k}}(x,\mathbf{z}/K)$. For the population size to be bounded by $K$, this imposes the following condition:
@@
Notice that for instance, classical epidemic models satisfy this condition, as the effective population size is often kept constant. Letting $Z^{K}=Z(X^{K})/K$ then ensures that $Z^{K}(t)\in\mathcal{Z}$ almost surely, for every $t\geq 0$.
Introducing the large population limit of the population process requires some notations. Consider a given sequence of initial conditions $(x,\mathbf{z}^{K})_{K\geq 1}$ such that for any $K$, $\mathbf{z}^{K}\in\mathcal{Z}_{K}/K$, $\mathbf{z}^{K}_{x}\geq 1/K$ and $\lim_{K\to\infty}\mathbf{z}^{K}=\mathbf{z}\in\mathcal{Z}$. Define $A(\mathbf{z})=(A_{x,y}(\mathbf{z}))_{x,y\in\mathcal{X}}$ for $\mathbf{z}\in\mathcal{Z}$ by $A_{x,y}(\mathbf{z})=\sum_{\mathbf{k}}(\mathbf{k}_{y}-1)\tau_{\mathbf{k}}(x,\mathbf{z})$. Throughout the section, we work under the following assumption.
@@
$\mathcal{G}f(x,\mathbf{z})=\sum_{\mathbf{k}}\tau_{\mathbf{k}}(x,\mathbf{z})\langle\mathbf{k}-e(x),f(\cdot,\mathbf{z})\rangle+\sum_{\mathbf{k},y}\mathbf{z}_{y}\tau_{\mathbf{k}}(y,\mathbf{z})\langle\mathbf{k}-e(y),\nabla_{\mathbf{z}}f(x,\mathbf{z})\rangle.$
@@
We are now ready to define the inhomogeneous spinal process. We want to sample at time $t>0$ in a population of initial condition $\mathbf{z}\in\mathcal{Z}$. At time $s\leq t$, the population is of composition $z(s)$, and the spine behaves as follows. Given its type $x$, it reproduces to leave descendance $\mathbf{k}$ and become of type $y$ at rate
@@
Equation (19) admits a unique strong solution $(\Upsilon^{(t)},z)$ in $\mathbb{D}([0,t],\mathcal{X})\times\mathcal{C}^{1}([0,t],\mathcal{Z})$. In addition, its semi-group $\mathcal{R}^{(t)}$ is a time-inhomogeneous semi-group of bounded linear operators on $\mathrm{L}^{\infty}(\mathcal{S})$, whose generator $\mathcal{A}^{(t)}$ is characterized by its action on functions $f\in\mathcal{C}^{1}(\mathcal{S})$ as follows. For $s\leq t$ and $(x,\mathbf{z})\in\mathcal{S}$,
@@
$\left|\mathbb{E}_{\mathbf{z}^{K}}\left[\sum_{u\in\mathbb{G}(t)}F((x_{u}(s),Z^{K}(s))_{s\leq t})\right]-\sum_{x\in\mathcal{X}}\mathbf{z}_{x}\mathfrak{m}(x,\mathbf{z},t)\mathbb{E}_{x,\mathbf{z}}\left[F((\Upsilon^{(t)}(s),z(s))_{s\leq t})\right]\right|\leq\frac{C}{\sqrt{K}}.$
This theorem shows that the spinal process $(\Upsilon^{(t)},z)$ indeed captures the typical lineage of a sampled particle in the large population limit $z$. Further, if one samples in the large population limit instead of a population whose size is of order $K$, the approximation error is of order $K^{-1/2}$, as expected for the large population limit.

### v3

@@ 1 Introduction
In order to address this limitation, there have been developments towards capturing the ancestral lineage of a sampled individual, as well as the whole genealogical tree, in populations with interactions. Recently, a spinal construction has been developed for this setting, focusing on multi-type processes with discrete type space [4]. The general idea consists in biasing the reproduction rates of the process, both along the spine and outside of it, according to a positive function $\psi$ of the reproducing particle’s trait $x$ and the population’s type composition $\mathbf{z}$. Intuitively, $\psi(x,\mathbf{z})$ can be regarded as the individual’s reproductive value or long-term fertility. Hence, when the spine reproduces, offspring with higher values of $\psi$ given the population state are favored, while the offspring of individuals outside of the spine are biased towards rendering the population more favorable for the spine. This spinal construction has since been extended to include more general type spaces [23]. It further has served to study the convergence of genealogies of density-dependent branching processes to the Kingman coalescent, establishing a connection with backward-in-time approaches [1].
@@
In this paper, we focus on the empirical type distribution of ancestral lineages ($\psi=1$), which arguably corresponds to the most natural and naive sampling strategy. In particular, we aim to capture the underlying survivorship bias. Notably, in the many-to-one formula associated to the aforementioned $\psi$-spine, spinal trajectories are penalized by an exponential weight which is generally stochastic. Thus the interpretation of the survivorship bias embodied by this spinal process is not straightforward, as this penalization needs to be taken into account. In addition, Monte-Carlo estimations of the many-to-one formula become delicate as rare trajectories may have a tremendous impact. As a consequence, we aim at proposing an alternative spinal construction, whose associated many-to-one formula does not require exponential weighting of trajectories. This is achieved by a time-inhomogeneous spinal process, inspired from a similar auxiliary process which captures the lineage of a uniformly sampled individual in a branching process with large initial population [22].
@@
Further, a natural regime to consider is the large population limit $K\to+\infty$. Under classical assumptions, several large population approximations can frequently be established. Indeed, these approximations may be more prone to mathematical analysis than the original population process. From a numerical point of view, simulating large population approximations is often less expensive than individual-based models for large population sizes, both in terms of computation time and memory. For instance, numerically evaluation of the $\psi$-spine many-to-one formula requires simulations of the spinal population process, which is computationally expensive for large population sizes. As a consequence, it is natural to develop approximate sampling strategies, which allow to sample directly in those large population approximations. Building on the large population approximation of the empirical distribution of ancestral lineages [4], we establish a many-to-one formula for sampling in the deterministic large population limit with our time-inhomogeneous spinal process. Finally, we quantify the approximation error which is made by sampling in the large population limit instead of the original population process.
@@ 3 Survivorship bias and the empirical distribution of ancestral lineages
In this section, our aim is to propose a many-to-one formula without stochastic penalization of spinal trajectories, in an effort to gain insight inon the survivorship bias. In order to do so, we will introduce a time inhomogeneous spinal construction. This approach is inspired by [22]. While we are particularly interested in the empirical distribution of ancestral lineages, the results obtained in this Section hold for a large class of sampling strategies, and are thus stated in a general setting.
@@ 3.1 A many-to-one formula
Let us now introduce the time-inhomogeneous spinal process, which allows to capture the behavior of the ancestral lineage of a $\psi$-weighted sample of the population process. For $t\geq 0$ fixed, we will consider the time-inhomogeneous Markov process $(Y^{(t)}(s),\zeta^{(t)}(s))_{s\leq t}$ defined as follows. The main idea is to follow the type $Y^{(t)}$ of a distinguished individual, which will be referred to as spine in analogy to classical spinal constructions. At time $s\leq t$, when of type $x$ in a population of state $\mathbf{z}$, the spine is replaced by offspring $\mathbf{k}$ and switches to type $y$ (i.e. among its offspring $\mathbf{k}$, one individual of type $y$ is chosen to become the new spine) with rate
@@
In other words, compared to the original process, at any time $s\leq t$, transitions along the distinguished lineage are biased in favor of those which lead to a larger $\psi$-average offspring at the final time $t$. However, due to the density-dependence of division rates, it is necessary to keep track of the population state $\zeta^{(t)}$. Again, transitions need to be biased, in order to account for the modified behavior of the distinguished individual when compared to the original process. As a consequence, when the population is in state $\mathbf{z}$ and the spine is of type $x$ at time $s\leq t$, individuals of type $y$ other than the spine is replaced by offspring $\mathbf{k}$ at rate
@@
Here, the bias favors those transitions which lead to a more favorable environment for the spine, i.e. a population composition in which the $\psi$-average of the spine’s offspring is high.
@@
$Y^{(t)}$ $(s)=Y^{(t)}(0)+\int_{0}^{s}\int_{E}\mathbf{1}_{\left\{\theta\leq\rho^{(t)}_{y,\mathbf{k}}(r,Y^{(t)}(r-),\zeta^{(t)}(r-))\right\}}(\mathbf{e}(y)-Y^{(t)}(r-))Q(dr,d\theta,dy,d\mathbf{k}),$ (5)
$\zeta^{(t)}$ $(s)=\zeta^{(t)}(0)+\int_{0}^{s}\int_{E}\mathbf{1}_{\left\{\theta\leq\rho^{(t)}_{y,\mathbf{k}}(r,Y^{(t)}(r-),\zeta^{(t)}(r-))\right\}}(\mathbf{k}-Y^{(t)}(r-))Q(dr,d\theta,dy,d\mathbf{k})$
$+\int_{0}^{s}\int_{E}\mathbf{1}_{\left\{\theta\leq(\zeta^{(t)}_{y}(r-)-\mathbf{1}_{\left\{Y^{(t)}(r-)=y\right\}})\widehat{\rho}^{(t)}_{\mathbf{k}}(r,y,Y^{(t)}(r-),\zeta^{(t)}(r-))\right\}}(\mathbf{k}-\mathbf{e}(y))\widehat{Q}(dr,d\theta,dy,d\mathbf{k}).$
@@
This result shows that the reproduction rates defined in Equations (3) and (4) yield the appropriate survivorship bias: reproduction events are favored if and only if they increase the $\psi$-average of the offspring at sampling time. In particular, the survivorship bias may thus be computed explicitly, giving access to a more precise interpretation. This will be illustrated in Section 3.3.
@@ 3.2.2 Proof of the many-to-one formula
Let us now turn to the inhomogeneous semi-group property. Let $r\leq\tau\leq s\leq t$, and consider $f:\mathcal{S}_{K}\to\mathbb{R}_{+}$ and $(x,\mathbf{z})\in\mathcal{S}_{K}$. Throughout the proof, we let $\mathfrak{X}_{0}=\mathfrak{X}(\mathbf{z})$. By definition of the semi-group,
@@
$\mathbb{E}[\sum_{\begin{subarray}{c}u\in\mathbb{G}(t)\\ u\succeq u_{x}(r)\end{subarray}}\psi(x_{u}(t),Z(t))m(x_{u}(\tau),Z(\tau),t-\tau)^{-1}G((x_{u}(\sigma),Z(\sigma))_{\sigma\leq\tau}))|X(r)=\mathfrak{X}_{0}]$ (15)
$=\mathbb{E}[\!\!\sum_{\begin{subarray}{c}v\in\mathbb{G}(\tau)\\ v\succeq u_{x}(r)\end{subarray}}\!\!\mathbb{E}[\!\!\sum_{\begin{subarray}{c}u\in\mathbb{G}(t)\\ u\succeq v\end{subarray}}\!\!\psi(x_{u}(t),Z(t))|X(\tau)]m(x_{v}(\tau),Z(\tau),t-\tau)^{-1}G((x_{v}(\sigma),Z(\sigma))_{\sigma\leq\tau}))|X(r)=\mathfrak{X}_{0}]$
$=\mathbb{E}[\sum_{\begin{subarray}{c}v\in\mathbb{G}(\tau)\\ v\succeq u_{x}(r)\end{subarray}}G((x_{v}(\sigma),Z(\sigma))_{\sigma\leq\tau}))|X(r)=\mathfrak{X}_{0}].$
@@ 3.3 An application to the empirical distribution of ancestral lineages in a density-dependent population
The particles behave as follows. If the population corresponds to a single particle of type $A$, the latter may give birth to another particle of the same type at rate $b$. Whenever there are two particles of type $A$, each may die due to competition at rate $a$, or escape competition by switching to state $B$ at rate $c_{A}$. Finally, a particle of type $B$ can only switch back to state $A$, at rate $c_{B}$. Throughout the following, we assume that $b=a$ and $c_{A}=c_{B}$, for ease of computation. In other words, the state space of the population with a distinguished particle is described by the ordered set $\{(A,1,0),(A,2,0),(A,1,1),(B,1,1)\}$, and the dynamics are characterized by four reproduction rates :
@@
$m(t)=\frac{4}{5}\begin{pmatrix}2\\ 1\\ 1\\ 1\\ \end{pmatrix}-\frac{\lambda_{-}}{10\sqrt{\Delta}}\mathrm{e}^{-\frac{\lambda_{+}t}{2}}\begin{pmatrix}-\frac{-3b+3c-\sqrt{\Delta}}{2c}\\ -\frac{-3b+c+\sqrt{\Delta}}{2c}\\ 1\\ 1\\ \end{pmatrix}+\frac{\lambda_{+}}{10\sqrt{\Delta}}\mathrm{e}^{-\frac{\lambda_{-}t}{2}}\begin{pmatrix}-\frac{-3b+3c+\sqrt{\Delta}}{2c}\\ -\frac{-3b+c-\sqrt{\Delta}}{2c}\\ 1\\ 1\\ \end{pmatrix}$,
@@ 4 Empirical distribution of ancestral lineages in large populations
In this section, we are interested in empirical distribution of ancestral lineages in large populations, corresponding to the case $\psi=1$. Indeed, when the population size $K$ grows to infinity, the population trait distribution converges under classical assumptions to a deterministic limit $z$, characterized as the unique solution of a dynamical system. Here, we show that it is possible to sample in the limit $z$ with a time-inhomogeneous spine construction, analogously to Theorem 3.3. In addition, we quantify the approximation error made by sampling in the large population limit instead of the original finite-size population process.
@@ 4.1 Many-to-one formula and speed of convergence
For $K\geq 1$, consider the population process $X^{K}$ where an individual of type $x$ in a population of composition $\mathbf{z}\in\mathcal{Z}_{K}$ is replaced by offspring $\mathbf{k}$ at rate $\tau_{\mathbf{k}}(x,\mathbf{z}/K)$. For the population size to be bounded by $K$, this imposes the following condition:
@@
Notice that for instance, classical epidemic models satisfy this condition, as population size is often kept constant. Letting $Z^{K}=Z(X^{K})/K$ then ensures that $Z^{K}(t)\in\mathcal{Z}$ almost surely, for every $t\geq 0$.
Introducing the large population limit of the population process requires some notations. Consider a given sequence of initial conditions $(x,\mathbf{z}^{K})_{K\geq 1}$ such that for any $K$, $\mathbf{z}^{K}\in\mathcal{Z}_{K}/K$, $\mathbf{z}^{K}_{x}\geq 1/K$ and $\lim_{K\to\infty}\mathbf{z}^{K}=\mathbf{z}\in\mathcal{Z}$. Define $A(\mathbf{z})=(A_{x,y}(\mathbf{z}))_{x,y\in\mathcal{X}}$ for $\mathbf{z}\in\mathcal{Z}$ by $A_{x,y}(\mathbf{z})=\sum_{\mathbf{k}}(\mathbf{k}_{y}-\mathbf{1}_{\left\{x=y\right\}})\tau_{\mathbf{k}}(x,\mathbf{z})$. Throughout the section, we work under the following assumption.
@@
$\mathcal{G}f(x,\mathbf{z})=\sum_{\mathbf{k}}\tau_{\mathbf{k}}(x,\mathbf{z})\langle\mathbf{k}-\mathbf{e}(x),f(\cdot,\mathbf{z})\rangle+\sum_{\mathbf{k},y}\mathbf{z}_{y}\tau_{\mathbf{k}}(y,\mathbf{z})\langle\mathbf{k}-\mathbf{e}(y),\nabla_{\mathbf{z}}f(x,\mathbf{z})\rangle.$
@@
We are now ready to define the inhomogeneous spinal process. We want to sample at time $t>0$ in a population of initial condition $\mathbf{z}\in\mathcal{Z}$. At time $s\leq t$, the population is of composition $z(s)$, and the spine behaves as follows. Given its type $x$, it is replaced by offspring $\mathbf{k}$ and become of type $y$ at rate
@@
Consider Assumptions 4.1 and 4.2. Equation (19) admits a unique strong solution $(\Upsilon^{(t)},z)$ in $\mathbb{D}([0,t],\mathcal{X})\times\mathcal{C}^{1}([0,t],\mathcal{Z})$. In addition, its semi-group $\mathcal{R}^{(t)}$ is a time-inhomogeneous semi-group of bounded linear operators on $\mathrm{L}^{\infty}(\mathcal{S})$, whose generator $\mathcal{A}^{(t)}$ is characterized by its action on functions $f\in\mathcal{C}^{1}(\mathcal{S})$ as follows. For $s\leq t$ and $(x,\mathbf{z})\in\mathcal{S}$,
@@
$\left|\frac{1}{K}E_{\mathbf{z}^{K}}\left[\sum_{u\in\mathbb{G}(t)}F((x_{u}(s),Z^{K}(s))_{s\leq t})\right]-\sum_{x\in\mathcal{X}}\mathbf{z}_{x}\mathfrak{m}(x,\mathbf{z},t)\mathbb{E}_{x,\mathbf{z}}\left[F((\Upsilon^{(t)}(s),z(s))_{s\leq t})\right]\right|\leq\frac{C}{K^{1/4}}.$
This theorem shows that the spinal process $(\Upsilon^{(t)},z)$ indeed captures the typical lineage of a sampled particle in the large population limit $z$. Further, if one samples in the large population limit instead of a population whose size is of order $K$, the approximation error is at most of order $K^{-1/4}$.